OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (amistre64):

y=2x-2(5)+3 maybe

OpenStudy (mani_jha):

Do you know the two-point form of a line? y-y1=m(x-x1) where (x1,y1) is a known point on the line. m=2 because the slope of the two lines should be same because of being parallel. Use the two-point form of an equation, y-y1=m(x-x1) y-5=2(x-3) y-5=2x-6 y-2x+1=0;

OpenStudy (amistre64):

except for, as is usually the case, a person transposes x for y

OpenStudy (amistre64):

(3,5) not= (5,3)